OpenStudy (anonymous):

whats the greatest common multiple of 3 and 4

OpenStudy (anonymous):

Don't you mean the smallest?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

3*4=12

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

12 as 3*4 = 12 and 4*3 = 12